Suppose that demand data for widgets made by XYZ Inc. is given below. Using an appropriate method (level , chase or mixed), plan the production (i.e. specify the production quantities and complete the table below) such that the total cost is made as small as possible. Regular time production capacity is 7000 units per month. There is no other method of production and the cost of regular time work force is fixed and it does not depend on the amount produced and hence may be omitted in the cost calculations. Holding cost per unit per month is $1.00, and shortage cost is $2.00 per unit per month, (short units must be supplied at a later period, but within the 4 month planning horizon.) what is the total inventory holding cost (use ending inventory to calculate this cost), total shortage cost and the total cost of your plan? September October November December Demand 4000 10000 8000 1000 Beginning Inventory 0 Production Shipments Shortage Ending Inventory Should be 1000 Shortage cost Inventory holding cost Total cost of the Plan:
Answer #1

Using level production strategy

Month Demand Production Inventory Shortage
0
September 4000 5750 1750 0
October 10000 5750 0 2500
November 8000 5750 0 4750
December 1000 5750 0 0
Total 1750 7250

total inventory holding cost = 1750*1 = 1750

Total shortage cost = 7250*2 = 14500

Total cost of the plan = total inventory holding cost +Total shortage cost = 1750+14500 = 16250